OpenStudy (anonymous):

suppose you had a dollar in your bank account.you spend $12 but have at least $51 left.how much money did you have initially?write and solve an inequality that represents this situation

OpenStudy (anonymous):

can someone plzs help me solve this

OpenStudy (♪chibiterasu):

Well basically we are removing 12 from the initial amount and the new total will be 51 so our first equation is: P - 12 >= 51 We use greater than or equal to sign since there is AT LEAST 51 left. To solve we just add 12 on both sides of the equation: P - 12 + 12>= 51 + 12 P >= 63 So the initial price is greater than or equal to $63

OpenStudy (anonymous):

ok thnks

OpenStudy (♪chibiterasu):

Your welcome :3